# How Global Calibration Strengthens Multiaccuracy

### v1

@@ A.1 Multiaccuracy as a Query Oracle
where $\chi_{S}(x)=\prod_{i\in S}x_{i}$ for $x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$ and $\widehat{f}(S)=\operatorname*{\mathbf{E}}_{x\sim\{-1,1\}^{n}}[f(x)\chi_{S}]$ denotes the Fourier coefficient of $f$ on $S$. Again defining the inner product $\langle f,g\rangle$ between two functions $f,g:\{-1,1\}^{n}\rightarrow\mathbb{R}$ as $\operatorname*{\mathbf{E}}_{x\sim\{-1,1\}^{n}}[f(x)g(x)]$ we see that the $2^{n}$ parity functions $\chi_{S}:\{-1,1\}^{n}\rightarrow\{-1,1\}$ form an orthonormal basis for the vector space of functions $\{-1,1\}^{n}\rightarrow\mathbb{R}$ [40, Thm. 1.5]. The Goldreich-Levin algorithm finds the largest Fourier coefficients of a function given membership access to it:
@@
Using Theorem A.3, we can observe the following simple corollary which shows that finding $p$ that is $(\mathcal{C},\tau)$-accurate for $\mathcal{D}$, where the marginal distribution over $\mathcal{X}$ is uniform, is at least as hard as Learning Parities with Noise (LPN). Since LPN is widely believed to be hard, we can conclude that the existence of an efficient algorithm that produces a $p$ that is $(\mathcal{C},\tau)$-multi-accurate for small values of $\tau$ is unlikely to exist, at least when the marginal distribution over $\mathcal{X}$ is uniform.
@@
When $\mathcal{C}$ is the class of parities, and $\mathcal{D}$ is such that the marginal distribution over $\mathcal{X}$ is uniform, then finding $p$ that is $(\mathcal{C},\tau)$-accurate for $\mathcal{D}$ and $\tau\leq 0.1$ is at least as hard as properly agnostically learning $\mathcal{C}$ with random classification noise with noise rate $0.1$.
@@
We know that the class of parities is efficiently properly agnostically learnable using the Goldreich-Levin or the Kushilevitz-Mansour algorithm[20, 33]. Since the noise rate is $0.1$, we know that there the optimal error is at most $0.1$, then using Theorem A.3 we can find a parity that has error at most $20.4+\varepsilon<1/$ 2. However, in the random classification noise setting all parities, except the target parity are uncorrelated with the target labels, so in fact we must have identified the target parity. ∎

### v2

@@ A.1 Multiaccuracy as a Query Oracle
where $\chi_{S}(x)=\prod_{i\in S}x_{i}$ for $x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$ and $\widehat{f}(S)=\operatorname*{\mathbf{E}}_{x\sim\{-1,1\}^{n}}[f(x)\chi_{S}]$ denotes the Fourier coefficient of $f$ on $S$. Again defining the inner product $\langle f,g\rangle$ between two functions $f,g:\{-1,1\}^{n}\rightarrow\mathbb{R}$ as $\operatorname*{\mathbf{E}}_{x\sim\{-1,1\}^{n}}[f(x)g(x)]$ we see that the $2^{n}$ parity functions $\chi_{S}:\{-1,1\}^{n}\rightarrow\{-1,1\}$ form an orthonormal basis for the vector space of functions $\{-1,1\}^{n}\rightarrow\mathbb{R}$ [41, Thm. 1.5]. The Goldreich-Levin algorithm finds the largest Fourier coefficients of a function given membership access to it:
@@
Using Theorem A.3, we can observe the following simple corollary which shows that finding $p$ that is $(\mathcal{C},\tau)$-accurate for $\mathcal{D}$, where the marginal distribution over $\mathcal{X}$ is uniform, is at least as hard as Learning Parities with Noise (LPN). Since LPN is widely believed to be hard, we can conclude that the existence of an efficient algorithm that produces a $p$ that is $(\mathcal{C},\tau)$-multiaccurate for small values of $\tau$ is unlikely to exist, at least when the marginal distribution over $\mathcal{X}$ is uniform.
@@
When $\mathcal{C}$ is the class of parities, and $\mathcal{D}$ is such that the marginal distribution over $\mathcal{X}$ is uniform, then finding $p$ that is $(\mathcal{C},\tau)$-multiaccurate for $\mathcal{D}$ and $\tau\leq 0.1$ is at least as hard as properly agnostically learning $\mathcal{C}$ with random classification noise with noise rate $0.1$.
@@
We know that the class of parities is efficiently properly agnostically learnable using the Goldreich-Levin or the Kushilevitz-Mansour algorithm [16, 37]. Since the noise rate is $0.1$, we know that there the optimal error is at most $0.1$, then using Theorem A.3 we can find a parity that has error at most $0.4+\varepsilon<1/2$. However, in the random classification noise setting all parities, except the target parity are uncorrelated with the target labels, so in fact we must have identified the target parity. ∎